# Evaluating language models as risk scores

### v1

@@ 4 Empirical findings
We use folktexts to evaluate several recently released models together with their instruction-tuned counterparts: the Llama 3 models [63], including the 8B and the 70B versions, the Mistral 7B [64], the Mixtral 8x7B and 8x22B variants [65], the Yi 34B [66], and the Gemma [67] 2B and 7B variants. All LLM results are obtained via zero-shot prompting. For comparison, results are also shown for a logistic regression (LR) model, and a gradient boosted decision trees model (XGBoost). The XGBoost model [68] is generally regarded as the state-of-the-art in tabular data tasks [69, 70].
@@
In this section we focus on the ACSIncome prediction task, which is the default folktexts benchmarking task. It consists in predicting whether a person’s income is above or below \$50K from 10 demographic features, and closely emulates the popular UCI Adult prediction task [71]. The evaluation test set consists of 160K randomly selected samples from the 2018 Census data. A separate set of 1.5M samples is used to train the supervised LR and XGBoost models, LLMs are used as zero-shot classifiers without fine-tuning. Experiments were ran on a cluster with Nvidia-A100-80GB GPUs, consuming an approximate total of 500 GPU hours. Appendix A presents additional results and plots for the experiments analyzed in this section, including the study of alternate prediction tasks.
Model ECE $\downarrow$ Brier score $\downarrow$ AUC $\uparrow$ Accuracy $\uparrow$ $\tau=0.5$ Accuracy $\uparrow$ $\tau=t^{*}$ Score $\sigma$ st. dev.
Llama 3 70B (it) 0.27 0.27 0.86 0.69 0.78 0.42
Llama 3 70B 0.20 0.20 0.86 0.70 0.77 0.14
Llama 3 8B (it) 0.32 0.30 0.85 0.62 0.78 0.37
Llama 3 8B 0.25 0.26 0.81 0.38 0.69 0.05
Mixtral 8x22B (it) 0.21 0.22 0.85 0.76 0.75 0.48
Mixtral 8x22B 0.17 0.19 0.85 0.68 0.77 0.10
Mixtral 8x7B (it) 0.16 0.18 0.86 0.78 0.78 0.42
Mixtral 8x7B 0.17 0.21 0.83 0.65 0.74 0.06
Mistral 7B (it) 0.21 0.22 0.83 0.77 0.77 0.42
Mistral 7B 0.20 0.23 0.80 0.73 0.74 0.04
Yi 34B (chat) 0.19 0.19 0.86 0.72 0.74 0.34
Yi 34B 0.24 0.22 0.85 0.62 0.74 0.20
Gemma 7B (it) 0.61 0.59 0.84 0.37 0.75 0.06
Gemma 7B 0.24 0.27 0.76 0.37 0.68 0.04
Gemma 2B (it) 0.63 0.63 0.73 0.37 0.63 0.00
Gemma 2B 0.14 0.25 0.62 0.45 0.63 0.02
LR 0.03 0.18 0.79 0.74 - -
XGBoost 0.00 0.13 0.90 0.82 - -
@@
Table 1: Zero-shot LLM results on the ACSIncome benchmark task, together with results for LR and XGBoost baselines fitted on 1.5M samples. Column “Acc. $\tau=t^{*}$” shows accuracy results after fitting a model binarization threshold on $n=100$ train samples, while “Acc. $\tau=0.5$” shows standard accuracy using the highest likelihood answer. Table cells are colored using a continuous color map between the worst ( in orange) and best ( in cyan) results of each column, where only the top 10% and bottom 10% of the scale is colored. Models generally achieve high predictive signal (high AUC) but poor calibration (high ECE).
@@ 4.1 Benchmark results
First, we perform a comprehensive evaluation of risk scores output by LLMs along multiple metrics. Results are summarized in Table 1.
@@
We first compare LLMs to the supervised learning baselines. We observe that a majority of LLMs (all of size 8B or larger) outperform the linear model in terms of area under the AUC curve. However, LLMs are clearly far from matching the supervised baselines with respect to calibration: all language models achieve high calibration error (ECE), while baselines achieve near-perfect calibration. Due to this high miscalibration, models struggle to translate scores with high predictive signal into high accuracy. We recall that AUC is agnostic to calibration, while accuracy on the maximum likelihood answer is not. All of the Gemma models have worse than random accuracy, and most LLMs fail to surpass even the linear baseline on this metric. Only the instruction-tuned Mistral models (7B, 8x7B, and 8x22B) outperform the linear model in terms of accuracy (see accuracy with $\tau=0.5$).
To show-case potential accuracy gains from improved calibration, we add the “Acc. $\tau=t^{*}$” column, where the risk-score binarization threshold $\tau\in[0,1]$ is fitted on a small set of $n=100$ random training samples. As expected, LLMs with poor calibration but high AUC see steep accuracy improvements after re-calibrating the threshold; e.g., Gemma 7b (it) jumps from 0.37 to 0.75 accuracy. Inspecting accuracy alone would fail to reveal the predictive signal of the risk scores.
@@
Finally, we focus on comparing base models to their instruction-tuned counterparts, marked with ‘(it)’ or ‘(chat)’. A striking trend is visible across the board: instruction-tuning generally worsens calibration. At the same time, we see small but consistent improvements in AUC after instruction-tuning. Appendix A.1 presents results on the four additional prediction tasks. The same trend of instruction-tuning leading to worse calibration and higher AUC is broadly replicated. However, performance across different tasks is somewhat inconsistent: LLMs generally have stronger predictive signal than a supervised linear model on the income and travel-time prediction tasks, but consistently underperform the linear baseline on the address change (mobility) task.
@@
Figure 2: Calibration curves for base (left) and instruction-tuned (right) models of the Llama and Mistral families. Curves are computed using $10$ quantile-based score bins.
@@ 4.2 Score distribution
To get additional insights into the difference between base models and their instruction-tuned counterparts, we inspect their risk scores more closely.
Figure 2 shows calibration curves of base and instruction-tuned models of the Llama and Mistral families. Both base and instruction-tuned models have poor score calibration, but failure modes are entirely different: Base models output under-confident scores, while instruction-tuned models output over-confident scores. To quantify this result, we introduce a measure of risk score confidence bias similar to the ECE metric:
@@
where $M$ is the number of score bins, $B_{m}$ is the set of samples in score bin $m$, $\mathrm{Acc}(.)$ is the accuracy on a given set of samples, and $\mathrm{Conf}(.)$ is the confidence on a given set of samples measured as the mean risk score for the highest likelihood class. Figure 3 shows the risk score confidence bias results. The two miscalibration modes are evident: confidence bias is higher for instruction-tuned models and lower (or even negative) for base models. That is, instruction-tuned models are generally over-confident in their predictions, outputting higher scores than their accuracy would warrant, while no such trend is visible for base models. Section 4.3 further analyzes subgroup calibration and its implications to algorithmic fairness.
Figure 4 shows the risk score distribution for a variety of model pairs. The score distributions for base and instruct model variants are immediately distinguishable: base models consistently produce low-variance distributions centered around 0.5, while instruction-tuned variants often output scores near 0 or 1. Even for the two versions of Llama 3 70B, which achieve exactly the same AUC, the score distributions are widely different. On this model pair, no predictive performance is gained by instruction tuning, but answers of the instruction-tuned model have significantly higher (over-)confidence. In fact, while the base Llama 3 70B has an average of 0.07 under-confidence bias, the instruct variant produces risk scores on average 0.21 over-confident (see Figure 3). The same trend is visible on all model pairs, with Yi 34B showing the smallest difference between base and instruct variants. Note that the instruct Gemma models degenerate into predicting only positive outcomes with high confidence score, hence why they have the lowest accuracy, worst Brier score, and worst ECE.
Interestingly, in contrast to knowledge testing where high confidence is typically associated with better models, our evaluation reveals that even if the true outcome has high entropy, instruction-tuning polarizes score distributions.
@@
Figure 3: Risk score confidence bias for all LLMs on the ACSIncome task. Negative values indicate under-confident risk scores (overestimating uncertainty), while positive values indicate over-confident risk scores (underestimating uncertainty). For all but the Yi model pair, the instruct model shows higher over-confidence.
@@
Figure 4: Risk score distribution for base and instruction-tuned model pairs on the ACSIncome task. After instruction-tuning, models exhibit high confidence, but worse calibration in general. The XGBoost scores showcase a perfectly calibrated distribution ($\text{ECE}\approx 0.00$). Consult Table 1 for details.
Figure 5: Shift of score distribution with increasing evidence on the ACSIncome task, for the Mixtral 8x7B model (which achieved the best Brier score). Features are described in Table A5. Score distribution gets more discriminative as more evidence is added, successfully increasing scores’ predictive signal (AUC).
@@ 4.3 Subgroup calibration
Next, we evaluate risk score calibration across different subpopulations. Figure 6 shows calibration curves for two sets of models on the ACSIncome task, evaluated on three subgroups specified by the instantiation of the race attribute in the data. We pick the three race categories with the largest representation in the ACS data. Appendix A.2 presents further discussion on subgroup miscalibration and its potential causes and implications. Note that a positive prediction of $Y=1$ is arguably the advantageous outcome, as it corresponds to the high-income category (“Earns above \$50,000 per year”).
[…]
Figure 6: Calibration curves for the Yi 34B and the Mixtral 8x22B models, across different race subpopulations on the ACSIncome task. Left: Base models. Right: Instruction-tuned models. The instruction-tuned models shown are worse offenders in terms of subgroup-specific score bias. The Yi 34B models (top) overestimate income, but the Black population sees consistently lower overestimates. The Mixtral 8x22B models (bottom) both over- and underestimate income depending on the score value, and output lower scores for the Black population.
4.4 Varying degree of uncertainty
Finally, we consider the dependence of risk scores on the available evidence. For this study we use the Mixtral 8x7B model, which achieves the best (lowest) Brier score among evaluated models (reflecting both high accuracy and high calibration). Figure 5 shows score distributions with increasing evidence: starting with only 2 features, and iteratively adding 2 features at a time. This sequence demonstrates how unrealizable, underspecified prediction tasks differ from realizable prediction tasks. Predicting income based on an individual’s place of birth (POBP) and race (RAC1P) is naturally not possible to a high degree of accuracy, forcing any calibrated model to output lower-confidence risk scores. Indeed, both base and instruct variants correctly output lower confidence scores for the smaller feature sets (left plots) when compared with the larger feature sets (right plots). Appendix A.3 goes further in-depth on how score distributions change with varying data uncertainty.
In addition to providing insights into risk scores, varying individual features in the prediction task also provides insights into LLM feature importance. Appendix A.4 presents LLM feature importance results and discusses the main differences to traditional supervised learning model.
5 Conclusion
@@
We introduced folktexts, a software package that provides datasets and tools to evaluate risk scores produced by language models. We believe that folktexts strengthens the evaluation ecosystem in a direction that was previously underserved. Unlike most existing LLM benchmarks, the datasets we introduced have inherent outcome uncertainty, making them useful as a basis for systematic evaluation of uncertainty quantification in LLMs.
At present, our empirical findings show that LLM risk scores generally have strong predictive signal, but are wildly miscalibrated. Such models may be good for knowledge testing, but lack adequate indicators of uncertainty. Additionally, we show that most instruction-tuned models have worse calibration but slightly higher accuracy and AUC than their base-model counterparts. Crucially, we demonstrate how inspecting calibration curves and score distributions is necessary to understand differences across models, such as the differences between base and instruct model variants, which are obfuscated when using single calibration metrics.
Going forward, folktexts features a modular, easily extensible architecture. The package therefore makes it easy to evaluate different models, prompting strategies, and metrics. We define reproducible reference standards for evaluating LLMs, but encourage researchers to use what they believe is best for different applications. We envision that future package extensions will include other uncertainty quantification methods, such as confidence intervals, or conformal prediction methods.

### v3

@@ 4 Empirical findings
We use folktexts to evaluate several recently released models together with their instruction-tuned counterparts: the Llama 3 models [66], including the 8B and the 70B versions, the Mistral 7B [67], the Mixtral 8x7B and 8x22B variants [68], the Yi 34B [69], and the Gemma [70] 2B and 7B variants. We also evaluate GPT 4o mini [71] through the OpenAI API (note that no base model version is available). Instruction-tuned models are marked ‘(i.t.)’. For comparison, results are also shown for a logistic regression (LR) model, and a gradient boosted decision trees model (XGBoost). The XGBoost model [72] is generally regarded as the state-of-the-art in tabular data tasks [73].
@@
In this section we focus on the ACSIncome prediction task, which is the default folktexts benchmarking task. It consists in predicting whether a person’s income is above or below \$50K from 10 demographic features, and closely emulates the popular UCI Adult prediction task [74]. The evaluation test set consists of 160K randomly selected samples from the 2018 Census data. A separate set of 1.5M samples is used to train the supervised LR and XGBoost models, LLMs are used as zero-shot classifiers without fine-tuning. Both multiple-choice prompting and chat-style numeric prompting were used to obtain two separate risk score distributions for each model (as described in Section 3.2). We focus on analyzing multiple-choice prompting results, as it is arguably the standard in LLM benchmarking [19, 20, 21, 22, 23, 24, 39, 45, 46, 47, 48]. Appendix A presents additional results and plots for the experiments analyzed in this section, including a more in-depth look into numeric prompting results, as well as results on alternative prediction tasks. Experiments were ran on a cluster with Nvidia-A100-80GB GPUs, consuming an approximate total of 500 GPU hours.
Model Multiple-choice prompting Numeric risk prompting
ECE $\downarrow$ Brier score $\downarrow$ AUC $\uparrow$ Acc. $\uparrow$ ECE $\downarrow$ Brier score $\downarrow$ AUC $\uparrow$ Acc. $\uparrow$
GPT 4o mini (it) 0.24 0.24 0.85 0.74 0.05 0.16 0.83 0.78
Mixtral 8x22B (it) 0.21 0.22 0.85 0.76 0.11 0.17 0.84 0.77
Mixtral 8x22B 0.17 0.19 0.85 0.68 0.13 0.18 0.82 0.74
Llama 3 70B (it) 0.27 0.27 0.86 0.69 0.25 0.23 0.84 0.67
Llama 3 70B 0.20 0.20 0.86 0.70 0.27 0.24 0.82 0.54
Mixtral 8x7B (it) 0.16 0.18 0.86 0.78 0.10 0.17 0.84 0.76
Mixtral 8x7B 0.17 0.21 0.83 0.65 0.07 0.17 0.81 0.78
Yi 34B (it) 0.19 0.19 0.86 0.72 0.22 0.21 0.80 0.48
Yi 34B 0.25 0.22 0.85 0.62 0.15 0.19 0.83 0.61
Llama 3 8B (it) 0.32 0.30 0.85 0.62 0.23 0.23 0.81 0.67
Llama 3 8B 0.25 0.26 0.81 0.38 0.14 0.24 0.63 0.40
Mistral 7B (it) 0.21 0.22 0.83 0.77 0.16 0.19 0.83 0.70
Mistral 7B 0.20 0.23 0.80 0.73 0.36 0.32 0.75 0.49
Gemma 7B (it) 0.61 0.59 0.84 0.37 0.33 0.30 0.78 0.42
Gemma 7B 0.24 0.27 0.76 0.37 0.15 0.20 0.80 0.73
Gemma 2B (it) 0.63 0.63 0.73 0.37 0.28 0.31 0.50 0.37
Gemma 2B 0.14 0.25 0.62 0.45 0.37 0.37 0.50 0.63
LR 0.03 0.18 0.79 0.74 0.03 0.18 0.79 0.74
XGBoost 0.00 0.13 0.90 0.82 0.00 0.13 0.90 0.82
@@
Table 1: Zero-shot LLM results on the ACSIncome benchmark task, together with results for LR and XGBoost baselines fitted on 1.5M samples. Table cells are colored using a continuous color map between the worst ( in orange) and best ( in cyan) results of each column. Models generally achieve high predictive signal (high AUC) but poor calibration (high ECE). Numeric prompting leads to improved ECE but worse AUC.
@@ 4.1 Benchmark results
We perform a comprehensive evaluation of risk scores output by LLMs along multiple metrics. Results are summarized in Table 1.
Multiple-choice prompting.
@@
We observe that a majority of LLMs (all of size 8B or larger) outperform the linear model baseline (LR) in terms of predictive power (AUC). However, LLMs are clearly far from matching the supervised baselines with respect to calibration: all language models achieve very high calibration error (ECE), while baselines achieve near-perfect calibration. Due to this high miscalibration, models struggle to translate scores with high predictive signal into high accuracy. In fact, while most models achieve high AUC, most models struggle to surpass the supervised linear baseline in terms of accuracy. We recall that AUC is agnostic to calibration, while accuracy on the maximum likelihood answer is not. All of the Gemma models have worse than random accuracy, despite having clearly above random AUC (random would be 0.5). Only the instruction-tuned Mistral models (7B, 8x7B, and 8x22B) outperform the linear model in terms of accuracy. Interestingly, while larger models achieve higher AUC, calibration is not reliably improved by model size — differences across model families are more pronounced than across model sizes.
@@
Finally, we focus on comparing base models to their instruction-tuned counterparts, marked with ‘(it)’. A striking trend is visible across the board: instruction-tuning generally worsens calibration (higher ECE) when using multiple-choice prompting. At the same time, we generally see improvements in AUC and accuracy after instruction-tuning. Appendix A.2 presents results on the four additional prediction tasks. The same trend of instruction-tuning leading to worse calibration and higher AUC is broadly replicated. However, performance across different tasks is somewhat inconsistent: LLMs generally have stronger predictive signal than a supervised linear model on the income prediction and travel-time prediction tasks, but consistently underperform the linear baseline on the address change task (ACSMobility).
Numeric prompting.
We observe broad improvements in calibration (lower ECE) and Brier score loss across instruction-tuned models when using numeric prompting. Figure 2 shows the change in ECE between both prompting schemes. The trend is clear across most instruct models across all five benchmark tasks. The Yi 34B model is the exception, showing outlier results throughout the different experiments we conduct. On the other hand, results for base models are less conclusive (see Figure A1). At the same time, numeric prompting leads to small but consistent drops in predictive power of risk scores (AUC) on 4 out of 5 benchmark tasks, including ACSIncome. Figure A2 shows these changes visually for all tasks. Finally, we point out that the GPT 4o mini model produces a surprisingly well-calibrated risk score distribution for the income prediction task ($\text{ECE}=0.05$). In fact, for each benchmark task, at least one LLM is able to produce a remarkably well-calibrated score distribution (although a different model for different tasks). This promising result points to the fact that some small amount of data will likely always be needed to properly evaluate LLMs capabilities to model human population statistics, but such modelling can often be done with high degree of confidence and a good understanding on which outputs might be wrong.
Figure 2: Change in calibration error (ECE) of instruction-tuned models when using numeric risk prompting (orange circles) versus multiple-choice prompting (blue squares). Improvement/deterioration is represented by green/red arrows, respectively. An overwhelming majority of model/task pairs see calibration improvements.
@@
Figure 3: Calibration curves for base and instruction-tuned versions of the largest models studied, on the ACSIncome task. Curves are computed using $10$ quantile-based score bins. Risk scores were generated using multiple-choice-style prompting (left plots) or numeric chat-style prompting (right plots).
@@ 4.2 Score distribution
To get additional insights into the difference between base models and their instruction-tuned counterparts, we inspect their risk score distributions more closely.
Figure 3 shows calibration curves of the largest LLMs studied, for both base and instruct variants and for both prompting schemes. When using multiple-choice prompting (left-most plots of Fig. 3), both base and instruction-tuned models have poor score calibration, but failure modes are entirely different: Base models output under-confident scores, while instruction-tuned models output over-confident scores. To quantify this result, we introduce a measure of risk score confidence bias similar to the ECE metric:
@@
where $M$ is the number of score bins, $B_{m}$ is the set of samples in score bin $m$, $\mathrm{Acc}(.)$ is the accuracy on a given set of samples, and $\mathrm{Conf}(.)$ is the confidence on a given set of samples measured as the mean risk score for the highest likelihood class. Figure 4(a) shows the risk score confidence bias results. The two miscalibration modes are evident: confidence bias is higher for instruction-tuned models and lower (or even negative) for base models. That is, instruction-tuned models are generally over-confident in their predictions, outputting higher scores than their accuracy would warrant, while no such trend is visible for base models. On the other hand, when using numeric prompting, the two aforementioned failure modes are no longer evident, and the differences between base and instruction-tuned models are blurred (right-most plots of Fig. 3). In fact, Figure 4(b) shows a trend reversal when using numeric prompting: Base models now show a higher over-confidence in their risk scores, while instruction-tuned models show approximately neutral score bias.
(a) Multiple-choice prompting.
(b) Numeric prompting.
@@
Figure 4: Risk score confidence bias for all LLMs on the ACSIncome task. Negative values indicate under-confident risk scores (overestimating uncertainty), while positive values indicate over-confident risk scores (underestimating uncertainty). Instruction-tuned models are generally over-confident when using multiple-choice prompting (Fig. 4(a)), but this bias is considerably diminished when using numeric prompting (Fig. 4(b)).
Figure 5 shows the risk score distribution for a variety of model pairs, produced using multiple-choice prompting. The score distributions for base and instruct model variants are immediately distinguishable: base models consistently produce low-variance distributions centered around 0.5, while instruction-tuned variants often output scores near 0 or 1. The same trend is visible on all model pairs, with Yi 34B showing the smallest difference between base and instruct variants. The score distributions produced by base/i.t. model pairs are markedly different, even among base/i.t. pairs achieving the exact same AUC; e.g., Llama 3 70B and Mixtral 8x22B. For the largest Llama (70B) and largest Mistral (8x22B) models, no predictive performance is gained by instruction tuning, but answers of the instruction-tuned models have significantly higher (over-)confidence and worse calibration. In fact, while the base Llama 3 70B has an average of 0.07 under-confidence bias, the instruct variant produces risk scores on average 0.22 over-confident (see Figure 4(a)). Note that the instruction-tuned Gemma models degenerate into predicting only positive outcomes with high confidence score, hence why they have the lowest accuracy, worst Brier score, and worst ECE.
Crucially, our evaluation reveals a previously unreported shortcoming of using multiple-choice prompting with instruction-tuned models: instruction-tuning polarizes score distributions, even if the true outcome has high entropy. Standard realizable knowledge testing benchmarks can easily disguise this polarization phenomenon as improper quantification of model uncertainty. In fact, it seems evident that it is improper quantification of uncertainty in general, regardless of underlying uncertainty in the modelled distribution. The following subsection goes further in-depth on the influence of data uncertainty in score distribution. Appendix A.1 analyzes numeric prompting results.
@@
Figure 5: Risk score distribution for base and instruction-tuned model pairs on the ACSIncome task, using multiple-choice prompting. After instruction-tuning, models exhibit high confidence, but worse calibration in general. The XGBoost scores showcase a perfectly calibrated distribution ($\text{ECE}\approx 0.00$).
4.3 Varying degree of uncertainty
Next, we consider the dependence of multiple-choice risk scores on the available evidence. For this study we use the Mixtral 8x7B model, which achieves the best (lowest) Brier score among evaluated models (reflecting both high accuracy and high calibration). We compute income prediction risk scores with increasing evidence: starting with only 2 features, and iteratively adding 2 features at a time (see results in Figure A7). This sequence demonstrates how unrealizable, underspecified prediction tasks differ from realizable prediction tasks. Predicting income based on an individual’s place of birth (POBP) and race (RAC1P) is naturally not possible to a high degree of accuracy, forcing any calibrated model to output lower-confidence risk scores. Indeed, both base and instruct variants correctly output lower confidence scores for the smaller feature sets when compared with the larger feature sets (compare left-most to right-most plots of Fig. A7). However, instruction-tuning still leads to a clear polarization of risk score distribution, regardless of true data uncertainty: Score variance for base models is in range $\sigma\in[0.02,0.06]$, while for i.t. models it’s in range $\sigma\in[0.16,0.41]$. Appendix A.4 goes further in-depth on how score distributions change with varying data uncertainty.
In addition to providing insights into risk scores, varying individual features in the prediction task also provides insights into LLM feature importance. Appendix A.5 presents LLM feature importance results and discusses the main differences to traditional supervised learning models.
Figure 6: Calibration curves for the Mixtral 8x22B (top) and the Yi 34B (bottom) models, across different race subpopulations on the ACSIncome task. Left: Multiple-choice-style prompting. Right: Numeric or chat-style prompting. The models shown are the worst offenders in terms of subgroup-specific score bias. On average, Black individuals see a lower score ($x$ axis) for the same true probability of positive outcomes ($y$ axis).
@@ 4.4 Subgroup calibration
Finally, we evaluate risk score calibration across different subpopulations, such as typically done as part of a fairness audit. Figure 6 shows calibration curves for two sets of models on the ACSIncome task, evaluated on three subgroups specified by the instantiation of the race attribute in the data. We pick the three race categories with the largest representation in the ACS data. Note that a positive prediction of $Y=1$ is arguably the advantageous outcome, as it corresponds to the high-income category (“Earns above \$50,000 per year”). On the two models shown, samples belonging to the ‘Black’ population see consistently lower scores for the same positive label probability when compared to the ‘Asian’ or ‘White’ populations. The Mixtral 8x22B (it) and Yi 34B (it) models shown are the worst offenders, with an average risk score difference between Asian and Black groups of 0.17 and 0.13 respectively; i.e., a Black individual will receive on average a 0.17 or 0.13 lower score than an Asian individual for the same true probability of high-income $\mathbb{P}\{Y=1\}$. This poses a higher bar for Black individuals to get a positive prediction of “high-income”. Note that the remaining models studied do not show such striking differences in group-conditional calibration. In fact, this score bias can be reversed for some base models, overestimating scores from Black individuals compared with other subgroups (results for all models shown in Appendix A.3). Such differences in score calibrations arguably warrant a more in-depth analysis that escapes the scope of this paper. We raise concerns regarding subgroup miscalibration, which should caution practitioners against using such scores in consequential domains without a comprehensive fairness audit.
5 Discussion
@@
We introduced folktexts, a software package that provides datasets and tools to evaluate risk scores produced by language models. Unlike most existing LLM benchmarks, the datasets we introduced have inherent outcome uncertainty, making them useful as a basis for systematic evaluation of uncertainty quantification in LLMs. While uncertainty on realizable tasks reflects only model uncertainty (i.e., whether the model is aware of its lack of knowledge), uncertainty on unrealizable tasks is itself a type of knowledge over the underlying data distribution.
Our empirical findings show that LLM risk scores produced using standard multiple-choice Q&A generally have strong predictive signal, but are wildly miscalibrated. Such models may be good for knowledge testing, but lack adequate indicators of uncertainty, making them unsuitable for synthetic data generation. Instruction-tuned models generally have worse calibration but slightly higher accuracy and AUC than their base-model counterparts on multiple-choice answers. In fact, instruction-tuning leads to marked polarization of multiple-choice answer distribution, regardless of underlying true data uncertainty. This reveals a general inability of instruction-tuned LLMs to quantify uncertainty using multiple-choice Q&A. At the same time, verbally querying models for numeric probability estimates considerably improves calibration of instruction-tuned models, but at a small cost in AUC. Going forward, we envision that future package extensions will include other uncertainty quantification methods, such as confidence intervals, or conformal prediction methods.